OpenStudy (anonymous):

Find the slope of the line whose equation is 6x - 3y = 12. 2 -2 -4 correct?

OpenStudy (anonymous):

The slope would be 2 and it passes through the Y axis at -4 I assume that you rewrote it like it did to get 3y=6x-12 then y=2x-4

OpenStudy (a1234):

6x - 3y = 12 Put it in slope-intercept form to find the slope: 6x - 3y = 12 -3y = -6x + 12 y = 2x - 4 The slope is 2.